OpenStudy (anonymous):

How would you find the x intercepts of f(x)=(x+4)^2-3?

OpenStudy (anonymous):

Take f(x)=0 So, 0=(x+4)^2 -3 Take 3 to other side, 3=(x+4)^2 Take square root on both sides, \[\sqrt{3}=x+4\] \[\sqrt{3}-4=x\]

OpenStudy (anonymous):

For finding x-intercepts take F(x)=0 For finding y-intercepts take F(x)=(x+4)^2 -3 with x from (x+4) as 0 :)

OpenStudy (anonymous):

Why is the Precalc book telling me to take out -1 of the xs? that's what's been throwing me off..

OpenStudy (amistre64):

duplicate questions are a hassle .....

OpenStudy (anonymous):

@Jamroz: You will get correct answer this way as well whatever method suits you, use it.

OpenStudy (amistre64):

dont forget to take "both" sqrts

OpenStudy (anonymous):

what do you mean @amistre64?

OpenStudy (amistre64):

what squared makes 3? there are 2 answers

OpenStudy (amistre64):

x^2 = 3 has 2 results

OpenStudy (anonymous):

Yeah so there's \[\pm \sqrt{3} -4\]

OpenStudy (anonymous):

yeah...

OpenStudy (amistre64):

maybe a more intuitive answer is: x^2 = 4 (2)^2 = 4 (-2)^2 = 4 so x= 2 and -2

OpenStudy (anonymous):

and the vertex is (-4,-3)?

OpenStudy (amistre64):

yes, thats the vertex

OpenStudy (anonymous):

wasn't in class today, but given the homework. sorry for sounding stupid, thanks

OpenStudy (amistre64):

yw

OpenStudy (anonymous):

at least, not you stay with doubts